OpenStudy (anonymous):

If f(x) = −2x − 1 and g(x) = 6x − 1, find f[g(−2)]

OpenStudy (anonymous):

first find \(g(-2)=6\times (-2)-1=-12-1=-13\) then find \(f(-13)\)

OpenStudy (anonymous):

so the answer is f -13